Two cars leave a city and head in the same direction. After 6 hours, the faster car is 66 miles ahead of the slower car. The slower car has traveled 294 miles. Find the speeds of the two cars.

Answer by stanbon(75887)

Two cars leave a city and head in the same direction. After 6 hours, the faster car is 66 miles ahead of the slower car. The slower car has traveled 294 miles. Find the speeds of the two cars.
-----------------------
Faster Car DATA:
time = 6 hrs ; rate = x mph ; distance = 6x miles
-------------------
Equation:
distance = 294+66 = 360 miles
6x = 360
rate = x = 60 mph (Faster car rate)
------
Slower car rate = distance/time = 294/6 = 46 mph
==========================
Cheers,
Stan H.
==========================